The following financial information is available for Crane Corporation.

(in millions)

2020

2019

Average common stockholders’ equity $2,566 $2,607
Dividends declared for common stockholders 292 623
Dividends declared for preferred stockholders 55 55
Net income 502 564



Calculate the payout ratio and return on common stockholders’ equity for 2020 and 2019.
Answer #1

Calculate payout ratio

2020 2019
Payout ratio 292/447 = 65% 623/509 = 122%
Return on Common Stockholder's equity 447/2566 = 17% 509/2607 = 20%

Payout ratio = Dividend for Common Stockholder/(Net income-Preferred dividend)